OpenStudy (anonymous):

Which expression is equivalent to 4^1/3 times 4^1/6

OpenStudy (anonymous):

The Equation written better: \[4^{1/3}*4^{1/6}\] The Choices: \[A: 2\] \[B: 16\] \[C: \sqrt[3]{4}\] \[D: \sqrt[6]{16}\]

jimthompson5910 (jim_thompson5910):

The bases are the same, so you add the exponents.

jimthompson5910 (jim_thompson5910):

1/3 + 1/6 = ???

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 } + \frac{ 1 }{ 6 } = \frac{ 1 }{ 9 }\]

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you cannot add the denominators like that

OpenStudy (anonymous):

Oh okay that's only with multiplication where you don't have to change the denominators into a common denominator. Oops. Brain fart. So it's: \[\frac{ 1 }{ 3 } = \frac{ 2 }{ 6 }\] \[\frac{ 2 }{ 6 } + \frac{ 1 }{ 6 } = \frac{ 3 }{ 6 } =\frac{ 1 }{ 2 }\]

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

so "4^1/3 times 4^1/6" turns into 4^(1/2)

jimthompson5910 (jim_thompson5910):

anything to the 1/2 power is really taking the square root that means \[\Large 4^{1/2} = \sqrt{4}\]

jimthompson5910 (jim_thompson5910):

because I'm using the rule \[\Large x^{m/n} = \sqrt[n]{x^m}\]

OpenStudy (anonymous):

Which means that the answer is 2?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

np